# LookAhead: Preventing DeFi Attacks via Unveiling Adversarial Contracts

### v4

@@ 3.2.1. DeFi Attacks
There are many types of security breaches with complicated root causes that could lure attackers into exploiting a victim contract. In different attack scenarios, an adversary could use different techniques to conduct their exploitation. In this section, we present these attack methods, define key concepts involved in the attack process, and restrict LookAhead’s detection scope to a certain area of focus.
@@
    Attacking through adversarial contracts. This pattern allows the attacker to execute complex attacking logic in a single transaction, hence avoid half-way failures thanks to the atomic nature of a blockchain transaction. Using this method, an attacker first deploys a smart contract on-chain containing the core attack logic, then initiate the attack by calling the entry point function of the adversarial contract. Most DeFi attacks follow this pattern, hence we focus on addressing this type of attacks.
Scope of Detection. We specify three attack scenarios that fall outside of the scope this study. 1) Attacks that do not require contract deployments. Those resulting from private key leakage or attackers directly invoking public functions of the victim contracts using EOAs are outside the scope of our detection. 2) A small portion of attacks are completed directly through the constructor function at the time of contract deployment. Since the adversarial contracts used in these attacks already exhibit execution traces of the actual attack at the time of deployment, we consider them as adversarial transactions, which fall outside the scope of our detection. 3) We do not place our attention on detecting attacks such as rug pulls as they can only be launched by privileged users.
LookAhead focuses on unveiling attacks with intention to cause financial loss for DeFi protocols by exploiting vulnerabilities in on-chain smart contracts. To identify an attack, our goal is to detect adversarial contracts that contain the core attack logic. Other contracts, such as those used solely for concealing profits to avoid detection, fake token contracts to facilitate the main attack, and counterattack contracts used by front-running bots will not be included in our analysis.
@@ 3.2.2. Adversaries
Attacker Assumptions. Based on previous research findings (Zhou et al., 2023), our study makes the following assumption about $\mathcal{A}$:
@@
    $\mathcal{A}$ do not initiate attacks by batching $tx_{deploy}$ and $tx_{first}$ together in an atomic transaction during contract deployment, meaning that $\mathcal{A}$ splits an attack into two stages following the second pattern described in Figure 4;
@@
    $\mathcal{A}$ hard-code exploitation targets’ addresses in their malicious contracts whenever possible, and the victim’s address will emerge before the first attack transaction $tx_{first}$.
@@ 3.3. System Model
Binary Lifting. We do not use any source code information from the contracts, but only the bytecode deployed on the blockchain to ensure our model is capable of analyzing any smart contract on EVM blockchains. To do that, LookAhead uses the open-source binary lifting and analysis library Gigahorse (Grech et al., 2019) to decompile the contract bytecode and generate intermeidate representation (IR). We then run a program to visualize the Control Flow Graph (CFG) based on the IR.
PSCFT and Feature Extraction. With the basic CFG, we use a specialized feature extraction program to pull out features including but not limited to: token-related call information, flashloan usage information, etc. Finally, we convert the CFG into a condensed text form (PSCFT) and populate it with semantic information such as function names and external call target address labels.
Classifier Inference. Input features are properly tokenized and formatted before they run through our trained classifiers. These ML models are made up of a transformer, an optimal candidate classifier, and a meta classifier trained using supervised machine learning methods for binary classification. They are able to output either the classification result $label_{pred}\in\{0,1\}$ indicating whether the contract is adversary ($1$) or not (00), or they can produce a confidence score $0\%\leq p_{pred}\leq 100\%$ representing the possibility of the input contract being malicious.
Defense Mechanisms. After an ongoing attack has been identified by LookAhead, there are several defense techniques that can potentially be used by defenders and victim contract owners. 1) To defend against $\mathcal{A}^{\prime}$, a notification will be sent to the targets and they will be given time to react within their capabilities to prevent further damage such as sending an emergency pause transaction to their contracts. Moreover, if contract owners are willing to give us or any LookAhead system operator admin permissions to execute their shutdown methods on behalf of them, we can automatically react to possible attacks in an even more timely manner. 2) To counter $\mathcal{A}$, we plan to develop a service that offer a blacklist-like feature to DeFi protocol developers that allows them to block any transactions initiated from a continuously updating list of attacker contract addresses. If enough protocol users and developers incorporate our service as part of their access control pipeline, they can take full advantage of the real-time protection abilities of LookAhead at minimal to no cost to their operations.
@@ 4.1. Contract Collection
To perform supervised training of our classifier models, we collect a large set of past and recent smart contracts. For the positive (adversarial) dataset, we curated a collection of adversarial contracts used in past real-world DeFi incidents. For the negative (benign) dataset, we follow a methodical process to choose samples that are very unlikely to be related to any attacks.
@@ 4.1.1. Adversarial Contracts
To collect adversarial contracts, we gathered information on DeFi incidents from various publicly available sources, such as GitHub repositories documenting DeFi incidents (DeF, 2024; Lea, 2024), and alerts from security companies (blo, 2024; cer, 2024; SlowMist, [n. d.]). We conduct a meticulous review of these incidents to identify adversarial contracts used by attackers. Between April 2020 and July 2024, we collected 506 DeFi attacks. Of these, $29$ attacks ($5.7\%$) involved direct attacks from EOAs without using contracts, and $37$ attacks ($7.3\%$) were executed directly within the contract constructor. Both cases fall outside the scope of our detection. While attacks spanned multiple blockchains, the majority were on Ethereum and BSC. To simplify the study, we exclude 65 incidents happened on other chains. This left us with $375$ adversarial contracts: $142$ from Ethereum and $233$ from BSC. Following prior research (Zhou et al., 2023), we classify the attacks into three categories: Untrusted or unsafe calls ($13.3\%$), Access control mistakes ($14.4\%$), and Coding mistake ($72.3\%$). The first category includes issues like reentrancy, the second covers access control flaws, and the third involves coding mistakes such as arithmetic errors and logic absences.
@@ 4.1.2. Benign Contracts
To construct benign samples, we first collected a total of $796,437$ contracts deployed between June 2022 and June 2024 using Google’s BigQuery service. And we analyzed the transaction history of these contracts, focusing on the number of unique addresses interacted with the contracts from June 2022 to August 2024. The result showed that the adversarial contracts in our dataset often have very few unique users, with $355$ of them ($94.7\%$) interacting with only a single address, namely the attacker’s account, and a maximum of $6$. This low interaction count is consistent with the behavior of adversarial contracts, which typically feature access control and are designed for single-use only to attack specific victims. Based on this observation, we select contracts with at least $10$ unique interacting addresses as the determining factor for their benign nature. This filtering resulted in $210,643$ benign contracts.
@@ 4.2. Feature Selection
Based on empirical observations, we construct a selection of features for use in LookAhead’s ML model training. The set consists of features spanning over both the implementation and the deployment stages of a DeFi attack, aiming to capture any potential unexpected patterns of behaviour that belong to attackers. In this section, we explain our rationale behind the set construction.
@@ 4.2.1. Deployment Features
Based on this taxonomy, we conduct a statistical analysis on the fund source of our contract samples. For our benign samples, more than three-quarter of them have Safe sources, approximately one-tenth of their labels are Anonymous. In contrast, the majority of the adversarial contracts have Anonymous fund sources, with only a small proportion having Safe labels.
Transaction Data: $\{value,input\_data\_length,gas\_used\}$. Depending on the contract’s design, the contract creator has the ability to transfer native tokens to the contract upon deployment. The amount transferred is indicated by the $value$ attribute associated with the transaction. Adversaries typically do not perform such a transfer. The input data field of a deployment transaction holds the contract’s bytecode, we measure its length to account for the size of the contract. Additionally, since the gas consumption of the deployment transaction is proportional to the complexity of the contract, we combine these metrics to effectively measure the scale of the contract.
@@ 4.2.2. Implementation Features
Based on empirical observations, some of the characteristics that can be used to determine whether a contract is malicious or not include: implemented functions, internal and external function calls. To capture other hidden or intrinsic properties of a contract, we also design a special technique called Pruned Semantic-Control Flow Tokenization (PSCFT) that provides a condensed representation of the behaviour of smart contracts.
Implemented Functions: $\{func\_count,flashloan\_callback\_count\}$. Since adversarial contracts typically do not involve complex business logic, they often implement fewer functions. In our dataset, benign contracts frequently incorporate $\geq 20$ public functions, while adversarial ones tend to implement much fewer. We also calculate the number of flashloan callback functions and the proportion they take up over all public functions. This choice is motivated by the fact that many attack scenarios require attackers to have significant token holdings, which most adversaries do not naturally possess. Consequently, they resort to utilizing flashloan to temporarily borrow a substantial amount of tokens and execute attack logic within corresponding callback functions. Since the size of lending pools are limited, attackers often borrow tokens from multiple sources, resulting in multiple flashloan callbacks. In our dataset, more than 60% of adversarial contracts implement flashloan callbacks, while almost all benign contracts do not involve such callbacks.
Function Calls: $\{token\_call\_count,max\_token\_call\_count,avg\_token\_call\_count,delegate\_c\\ all\_count,selfdestruct\_count\}$. We take into consideration external calls related to ERC-20 tokens, including their count and proportion. Generally, the objective of DeFi attacks is to gain financial benefits, i.e., acquire valuable tokens. Therefore, external calls within adversarial contracts often involve token-related functions. The token-related functions we selected include, but are not limited to, common token standard (ERC, 2024a, b) functions and popular DEX (Uniswap, 2024; Adams et al., [n. d.]) functions. In our dataset, over $90\%$ of benign contracts do not have token-related calls. In contrast, almost all adversarial contracts involve token-related calls, with over two-thirds of them even having $10$ or more.
Furthermore, we introduce two additional features to assess the behaviour of public functions in a smart contract. $max\_token\_call\_count$ is designed to capture the maximum number of token-related interactions within the execution flow of any public function of a contract, which helps us identify contracts with exceptionally high token interactions. Moreover, $avg\_token\_call\_count$ is also measured to demonstrate the average number of token-related calls in contracts.
@@
Code Semantics: $\{PSCFT\}$. To understand the behaviours exhibited by DeFi attack adversaries and capture features that are potentially hidden from our empirical observations, we design a new representation for smart contracts called Pruned Semantic-Control Flow Tokenization ($PSCFT$). PSCFT improves upon the intermediate representation (IR) generated by Gigahorse to produce richer contextual information about the implementation of a contract. After constructing CFGs based on IR and going through an augmentation pipeline that retrieves and replaces function signatures and external addresses with appropriate semantic labels in the CFGs, PSCFT can effectively present a textual summary of function canonical names, control-flow graphs and external call chains. Figure 6.3 shows a contract PSCFT excerpt. One of the benefits of using PSCFT over IR is that IR alone does not provide useful information such as call target labels, and the size of the tokenization vocabulary is less manageable, rendering it not ideal for NLP training.
@@ 5.1. External Data Retrieval
Some of the features used in our dataset are obtained from external sources, we identify related third-party services in the following section. Based the data, we develop algorithms such as fund source tracing to retrieve specific fields to be used for both training and inference of our ML models.
Transaction Data. In this study, we focus on investigating and demonstrating the viability and effectiveness of the LookAhead system. Therefore, for the sake of simplicity, instead of running our own blockchain nodes, we obtain transaction raw data from an RPC (Remote Procedure Call) provider such as Alchemy (Alc, 2024) that relays our requests directly to other nodes in the blockchain network. Prior to the practical deployment of LookAhead, we recommend setting up dedicated nodes for minimized data processing delays. We store and format contract deployment transactions and their associated attributes such as nonce, value, and gas consumption, along with the sender and input data properties for use in subsequent phases of our feature extraction pipeline.
@@
Furthermore, we develop a fund source labeling algorithm that recursively follow the transaction flow and trace all the way back to the original fund holder. Utilizing the address label cloud (lab, 2024) and public name tag provided by chain explorers, we manually review and assign fund source labels to each group of addresses. As an example, Tornado.Cash was given the label of Anonymous, while centralized exchanges like Coinbase were given the label of Safe.

### v6

@@ 3.2.1. DeFi Attacks
There are many types of security breaches with complicated root causes, potentially leading to attacks on victims. In various scenarios, an adversary could use distinct techniques for exploitation. In this section, we present these attack methods, define key concepts involved in the attack process, and restrict LookAhead’s detection scope to a certain area of focus.
@@
    Attacking through adversarial contracts. This pattern allows the attacker to execute complex attacking logic in a single transaction, hence avoid half-way failures thanks to the atomic nature of a blockchain transaction. Using this method, an attacker first deploys a smart contract on-chain containing the core attack logic, then initiate the adversarial transaction by calling the entry point function of the adversarial contract. Most DeFi attacks follow this pattern, hence we focus on addressing this type of attacks.
Scope of Detection. We specify two attack scenarios that fall outside of the scope this study. 1) Attacks that do not require contract deployments. Those resulting from private key leakage or attackers directly invoking public functions of the victim contracts using EOAs are outside the scope of our detection. 2) A small portion of attacks are completed directly through the constructor function at the time of contract deployment. Since the adversarial contracts used in these attacks already exhibit execution traces of the actual attack at the time of deployment, we consider them as adversarial transactions, which fall outside the scope of our detection.
LookAhead focuses on unveiling attacks that intend to cause financial loss for DeFi protocols by exploiting vulnerabilities in on-chain smart contracts. To identify an attack, our solution detects adversarial contracts that contain the core attack logic. Other contracts, such as those used solely for concealing profits, fake token contracts that facilitate the main attack, and counterattack contracts used by front-running bots will not be included for our analysis.
@@ 3.2.2. Adversaries
Attacker Assumptions. Based on previous research findings (Zhou et al., 2023) (IEEE S&P’23), our study makes the following assumption about $\mathcal{A}$:
@@
    $\mathcal{A}$ does not initiate attacks by batching $tx_{deploy}$ and $tx_{first}$ together in an atomic transaction during contract deployment, meaning that $\mathcal{A}$ splits an attack into two stages following the second pattern described in Figure 2;
@@
    $\mathcal{A}^{\prime}$ hard-codes exploitation targets’ addresses in their adversarial contracts whenever possible, and the victim’s address will emerge before the first adversarial transaction $tx_{first}$.
@@ 3.3. System Model
Binary Lifting. We do not use any source code information from the contracts, but only the bytecode deployed on the blockchain to ensure our model is capable of analyzing any smart contract on EVM blockchains. To do that, LookAhead uses the binary lifting and analysis library Gigahorse (Grech et al., 2019) to decompile the contract bytecode and generate intermeidate representation (IR).
PSCFT and Feature Extraction. With the raw IR, we use a specialized feature extraction program to pull out features including but not limited to: token-related call information, flashloan usage information, etc. Finally, we convert the IR into a condensed text form (PSCFT) and populate it with semantic information such as function names and external call target address labels.
Classifier Inference. Input features are properly tokenized and formatted before they run through our trained classifiers. These ML models are made up of a transformer, an optimal candidate classifier, and a meta classifier trained using supervised machine learning methods for binary classification. They are able to output either the classification result $label_{pred}\in\{0,1\}$ indicating whether the contract is adversarial ($1$) or not (00), or they can produce a confidence score $0\%\leq p_{pred}\leq 100\%$ representing the possibility of the input contract being adversarial.
Defense Mechanisms. After detecting an ongoing attack, LookAhead enables several potential defense mechanisms for defenders and victim contract owners: 1) To defend against $\mathcal{A}^{\prime}$, notifications are sent to targets, allowing them to take defensive actions, such as executing an emergency pause. If contract owners grant admin permissions to LookAhead, the system can autonomously trigger shutdown methods on behalf of them. 2) To mitigate against $\mathcal{A}$, we propose a blacklist service for DeFi protocols, enabling them to block transactions from flagged adversarial contracts.
@@ 4.1. Contract Collection
To perform supervised training of our classifier models, we collect a large set of smart contracts from the Ethereum and BSC. For the positive (adversarial) dataset, we curated a collection of adversarial contracts used in real-world DeFi incidents. For the negative (benign) dataset, we follow a methodical process to collect contract samples that are very unlikely to be related to any attacks.
@@ 4.1.1. Adversarial Contracts
To collect adversarial contracts, we gather information on DeFi incidents from various publicly available sources, such as GitHub repositories documenting DeFi incidents (DeF, 2024; Lea, 2024), and alerts from security companies (blo, 2024; cer, 2024; slo, 2024). We conduct a meticulous review of these incidents to identify the contracts used by attackers. Between April 2020 and July 2024, we collected 506 DeFi attacks. In these, $29$ attacks ($5.7\%$) involved direct attacks from EOAs without using contracts, and $37$ attacks ($7.3\%$) were executed directly within the contract constructor. Both cases fall outside the scope of our detection. While attacks spanned multiple blockchains, the majority were on Ethereum and BSC. To simplify the study, we exclude 65 incidents happened on other chains. This left us with $375$ adversarial contracts: $142$ from Ethereum and $233$ from BSC. Following prior research (Zhou et al., 2023), we classify the attacks into three categories: 1) Untrusted or unsafe calls ($13.3\%$), 2) Access control mistakes ($14.4\%$), and 3) Coding mistake ($72.3\%$). The first category includes issues like reentrancy, the second covers access control flaws, and the third involves coding mistakes such as arithmetic errors and logic absences.
@@ 4.1.2. Benign Contracts
As contract-based DeFi attack detection approach is still a largely under-explored research direction, to date, there is no publicly available dataset that provides a comprehensive and representative set of ground truth benign contracts.
Consequently, we need to build our benign dataset from scratch. Given the vast number and variety of smart contracts deployed on the blockchain, constructing a large-scale dataset with sufficient sample diversity and label accuracy requires extensive manual review efforts and is prohibitively time-consuming. In response to this, previous ML-based work Forta (For, 2024) follow a heuristic approach for collecting benign samples.
Previous Approach. Forta assumes that all contracts verified on the chain explorer are benign. Despite being an intuitive design choice, analysis of our attack samples reveals that a small proportion of adversarial contracts are nonetheless verified. This may be because verification is an inexpensive task (often automated by development toolkits) and it has negligible impact on attack effectiveness, as existing detection methods do not focus on identifying adversarial contracts. Consequently, a more robust collection method is needed.
Key Observations. A smart contract can be deployed either 1) through a normal transaction, where its to_address field is NULL with bytecode included in the input_data field, or 2) via an internal transaction (also known as a message call) initiated by an existing contract during its execution. Between June 2022 and June 2024, the number of contracts created via internal transactions is significantly higher than those created via normal transactions (approximately 20 times more). A prior research (Gupta and Shukla, 2019) that analyzed $1.16$ million contracts created via internal transactions found that these contracts were generated by only $9,228$ creator contracts, indicating that a small number of contracts were responsible for deploying a vast number of contract instances. Most of these instances originate from factory contracts, e.g., Uniswap’s factory contract  [^2] https://etherscan.io/address/0x5c69bee701ef814a2b6a3edd4b1652cb9cc5aa6f, which are designed to repeatedly create similar contracts. Additionally, we observe that all adversarial samples in our dataset are deployed through normal transactions. These contracts are typically invoked by a very limited number of unique addresses. Specifically, $355$ of them ($94.7\%$) interacting with only a single address, namely the attacker’s account, and a maximum of $6$ unique users. This low interaction count is consistent with the behavior of adversarial contracts, which typically feature access control and are designed for single-use only to attack specific victims.
Our Method. Based on the observations, we construct our benign dataset following two key principles to ensure evaluation fairness, as well as dataset diversity and accuracy.
  • To ensure fair evaluation, we focus solely on contracts created via normal transactions, consistent with the deployment patterns of all adversarial samples, while avoiding internal transactions that could drastically inflate the dataset with numerous near-identical instances.
  • We select contracts with over a defined number of unique interacting addresses as an indicator of their benign nature, aiming to ensure reliability while maintaining sample diversity.
We first query Google’s BigQuery service to collect contracts deployed via normal transactions between June 2022 and June 2024, identifying $796,437$ contracts. Then we analyzed their transaction history, focusing on the number of unique interacting addresses from June 2022 to August 2024, and retained contracts with at least $10$ unique interacting addresses, resulting in a final benign dataset of $210,643$ contracts. While no heuristic method can fully eliminate potential bias, our approach offers a pragmatic trade-off between dataset scale, diversity, and correctness.
@@ 4.2. Feature Selection
Based on empirical observations, we construct a selection of features for use in LookAhead’s ML model training and inference. The set consists of features spanning both the implementation and the deployment stages of a DeFi attack, aiming to capture any potential unexpected patterns of behaviour exhibited by attackers. In this section, we explain our rationale behind the choice of features and provide supporting statistical data whenever applicable.
@@ 4.2.1. Deployment Features
Statistical Data: In benign samples, we observe that more than $75\%$ of them have Safe sources, and approximately $10\%$ of them are funded through Anonymous sources. In contrast, more than $70\%$ of the adversarial contracts have Anonymous sources, with only about 10% having Safe labels.
Transaction Data: $\{value,input\_data\_length,gas\_used\}$. Depending on the contract’s design, the contract creator can transfer native tokens to the contract upon deployment. Adversaries typically do not perform such a transfer. To avoid the potential discrepancy across Ethereum and BSC, we binarize the value attribute to indicate whether such a transfer occurred (true or false), rather than using the exact amount. The input data field of a deployment transaction holds the contract’s bytecode, we measure its length to account for the size of the contract. Additionally, we use $gas\_used$ to measure the computational resources consumed during deployment, which depends on the bytecode complexity and EVM execution, rather than gas fees, which relies not only on $gas\_used$ but also on the dynamically fluctuating gas price influenced by network congestion.
Statistical Data: In our dataset, about 80% of benign contracts have input data exceeding 15,000 bytes, while over 50% of adversarial contracts fall below this threshold. To eliminate potential EVM-induced differences in $gas\_used$ between Ethereum and BSC, we simulate the deployment of all BSC adversarial samples on Ethereum using identical bytecode to obtain a consistent gas_used value. However, 24% of the contracts failed due to interactions with contracts absent on Ethereum. Among the successfully samples, 94% showed less than a 1% deviation compared to original value. Given this negligible difference, for the failed samples, we directly use their original gas_used value.
@@
Statistical Data: We observe in our dataset that only $6$ adversarial contracts ($1.6\%$) are verified , while more than $85\%$ of benign contracts are verified.
@@ 4.2.2. Implementation Features
Based on empirical observations, some of the characteristics that can be used to determine whether a contract is adversarial or not include: implemented functions, internal and external function calls. To capture other hidden or intrinsic properties of a contract, we also design a special technique called Pruned Semantic-Control Flow Tokenization (PSCFT) that provides a condensed representation of the behaviour of smart contracts.
Implemented Functions: $\{func\_count,flashloan\_callback\_count\}$. Since adversarial contracts typically do not involve complex business logic, they often implement fewer functions. We also calculate the number of flashloan callback functions and the proportion they take up over all public functions. This choice is motivated by the fact that many attack scenarios require attackers to have significant token holdings, which most adversaries do not naturally possess. Consequently, they resort to utilizing flashloan to temporarily borrow a substantial amount of tokens and execute attack logic within corresponding callback functions. Since the size of lending pools are limited, attackers may borrow tokens from multiple sources, resulting in multiple flashloan callbacks.
Statistical Data: We can observe from the dataset that over 75% of benign contracts incorporate at least 20 public functions, whereas more than 75% of adversarial contracts implement 10 or fewer. Additionally, over 60% of adversarial contracts include flashloan callbacks, while more than 99% benign contracts involve no such callbacks.
Function Calls: $\{token\_call\_count,max\_token\_call\_count,avg\_token\_call\_count,delegate\_c\\ all\_count,selfdestruct\_count\}$. We take into consideration external calls related to tokens, including their count and proportion. Generally, the objective of DeFi attacks is to gain financial benefits, i.e., acquire valuable tokens. Therefore, external calls within adversarial contracts often involve token-related functions. The token-related functions we selected include, but are not limited to, common token standard (ERC, 2024a, b) functions and popular DEX (Adams et al., 2021) functions.
Furthermore, we introduce two additional features to assess the behaviour of public functions in a smart contract. $max\_token\_call\_count$ is designed to capture the maximum number of token-related interactions within the execution flow of any public function of a contract, which helps us identify contracts with exceptionally high token interactions. Moreover, $avg\_token\_call\_count$ is also measured to demonstrate the average number of token-related calls per public function.
@@
Statistical Data: It can be observed from our dataset that over $90\%$ of benign contracts do not have token-related calls. In contrast, almost all adversarial contracts involve token-related calls, with over $70\%$ of them having $10$ or more.
Figure 4. Side-by-side comparison of raw IR and PSCFT.
Code Semantics: $\{PSCFT\}$. To understand the behaviours exhibited by DeFi attack adversaries and capture features that are potentially hidden from our empirical observations, we design a new representation for smart contracts called Pruned Semantic-Control Flow Tokenization ($PSCFT$).
In smart contracts, functionalities (including potential attack logic) are implemented through functions, and interactions with external entities, such as invoking the functions implemented in other deployed contracts, are performed via message calls. For adversarial contracts, such interactions (e.g., with victim contracts) are indispensable. Therefore, each function, along with its message calls and control flows, is capable of capturing the intrinsic behavioral characteristics of a contract, providing key information for determining whether the contract is adversarial or benign.
To extract these critical features, we use Gigahorse (Grech et al., 2019) to lift the contract bytecode into an intermediate representation (IR), followed by a pruning process that remove extraneous elements while reconstructing control flows. To provide richer contextual information, the refined IR is fed into an augmentation pipeline that retrieves and replaces function signatures and external addresses with appropriate semantic labels. The resulting output is the fully PSCFT, which effectively represents a textual summary of function canonical names, control-flow graphs and external call chains. Figure 4 presents an excerpt comparing the raw IR of a contract with the corresponding PSCFT. We provide the detailed construction process of PSCFT in §5.2.
@@ 5.1. External Data Retrieval
Some of the features used in our dataset are obtained from external sources, we identify related third-party services in this section. Based the data, we use a algorithm to determine fund source for the contracts, enabling subsequent training and inference of our machine learning models.
Transaction Data. In this study, we focus on investigating and demonstrating the viability and effectiveness of the LookAhead system. Therefore, for the sake of simplicity, instead of running our own blockchain nodes, we obtain transaction raw data from an RPC (Remote Procedure Call) provider such as Alchemy (Alc, 2024) that relays our requests directly to other nodes in the blockchain network. Prior to the practical deployment of LookAhead, we recommend setting up dedicated nodes for minimized data processing delays. We store and format contract deployment transactions and their associated attributes such as nonce, value, and gas used, along with the sender and input data properties for use in subsequent phases of our feature extraction pipeline.
Figure 5. LookAhead’s system building workflow
@@
Furthermore, we develop a fund source labeling algorithm to determine a contract’s fund source. We first construct an address label dataset using the address label cloud (lab, 2024b, a) and the public name tags provided by chain explorers, following the definitions in §4.2.1. As an example, Tornado.Cash (Tor, 2024) is given the label of Anonymous. The algorithm traces fund flows from the contract deployer, recursively identifying the earliest funding source. If the source address has a known label, it is assigned; otherwise, tracing continues up to a predefined threshold depth. If no label is identified within this limit, the fund source of the contract is marked as Unknown.